# Patch of Invisibility: Naturalistic Black-Box Adversarial Attacks on Object Detectors

### v5

@@ Abstract
Adversarial attacks on deep learning models have received increased attention in recent years. Work in this area has mostly focused on gradient-based techniques, so-called “white-box” attacks, where the attacker has access to the targeted model’s internal parameters; such an assumption is usually untenable in the real world. Additionally, some attacks use the entire pixel space to fool a given model, which is neither practical nor physical. To accommodate these problems we propose the BBNP algorithm (Black-Box Naturalistic Patch): a direct, black-box, naturalistic, gradient-free method that uses the learned image manifold of a pretrained, generative adversarial network (GAN) to generate naturalistic adversarial patches for object detectors. This method performs model-agnostic black-box naturalistic attacks on object detection models by relying solely on the outputs of the model. Comparing our approach against five models, five black-box and two white-box attacks, we show that our proposed method achieves state-of-the-art results, outperforming all other tested black-box approaches. The code is available on GitHub at https://github.com/razla/Patch-of-Invisibility.
@@ 3 Method
We exploit the smaller dimensionality of the GAN’s latent space by using an Evolution Strategy algorithm [30]. This method updates the latent vector z iteratively based on feedback from the fitness function. To compare, optimizing a patch directly in pixel space (of size $3\times 256\times 256$, as in our experiments) would involve $196,608$ pixels. In contrast, our approach operates in a $128$-dimensional latent space, achieving a reduction in dimensionality by a factor of $1{,}000$. This reduction makes the optimization process faster and the generated attacks more natural, thanks to the pretrained GAN, which is trained on ImageNet [31].
@@
Figure 1: Naturalistic Black-Box Adversarial Attack: Overview of framework. The system creates patches for object detectors by using the learned image manifold of a pretrained GAN generator ($G$) on real-world images. We use a pretrained classifier ($C$) to force the optimizer to find a patch that resembles a specific class, the $TV$ component, which refers to total variation, aims to ensure the images are rendered with maximal smoothness, and the detector ($D$) for the actual detection loss. Efficient sampling of the GAN images via an iterative evolution strategy ultimately generates the final patch. The snowflake icon means frozen model weights.
@@ 3.2 Adversarial gradient estimation
$t$ is the timestep (epoch), $\alpha$ is the step size, and ~$\tilde{\nabla}_{z}\mathcal{L}_{total}$ is the gradient approximation of the total loss with respect to the latent vector $z$. We used $\tau=10$ in all experiments based on qualitative observations as follows. Figure 2 displays the patches generated without any form of projection. Conversely, Figure 5 presents the patches generated through the application of Equation 6.
@@ 4.1 Experimental setup
The experiments were performed using the Adam [38] optimizer, with a learning rate of $0.02$ and $\beta_{1}=0.5,\beta_{2}=0.999$. The optimization process involved populating the gradients through our evolutionary algorithm. If the changes in losses remain below $41e-$ 4 for at least 50 epochs, we reduce the learning rate, following [10]. Our generator consists of BigGAN2 [27] (with frozen weights), which has been pretrained on the ImageNet dataset, with a latent vector of size $120$, and $128\times 128$ output resolution. As the BigGAN generator is class-conditional, the generated patch can be guaranteed to belong to a specific class. We used five different object detectors—Tiny-YOLOv3 [39], Tiny-YOLOv4 [40], YOLOv5s [41] [15, 42], SSD-L [43, 44] and L-DETR [45] trained on the COCO dataset [46]—with an input resolution of $416\times 416$. We chose these faster, memory-efficient models so as to be able to conduct more experiments. Moreover, these models represent architectures that are predominantly deployed on edge hardware and widely utilized in production environments [47, 48, 49], thereby enhancing the practical relevance and realism of the experiments conducted. For each of the detected bounding boxes, we placed an adversarial patch of size $25\%$ of the bounding box. We adopt the method proposed by Hu et al. [10] and generate a mask encompassing the clothing region for each individual, onto which we position our adversarial patch.
@@
We compared our proposed attack to five other black-box methods and two white-box methods: Square Attack [35], which exploits the CNN vulnerability to square perturbations; Sparse-RS [34], which leverages [35] to craft minimal modifications of squares; RS (pixel space random search) [34], which uses random search on pixel space; GAN Latent RS (latent-space random search) which uses random search in GAN latent space; and NES (natural evolution strategies) [36], which is an evolutionary attack performed in pixel space. We employed a consistent query budget of 110,000 for all attacks, derived from our chosen population size of 110 and 1,000 epochs.  [^1] We slightly adjusted some of the compared methods to accommodate the object-detection task: We removed the limitation on $\epsilon$ to encourage broader exploration during optimization in competitive attacks, because perturbation limits are not part of our targeted scenario, which is patch attacks. In our implementation of Square Attack, Sparse-RS, and the two white-box attacks, WB-Adv and WB-Adv as proposed by Thys et al. [6] and Hu et al. [10], we adhered strictly to the default parameters specified in their respective works. For NES, RS, and Latent-RS, we employed a noise standard deviation of $\sigma=0.1$. Additionally, for Latent-RS, we utilized two distinct classifier guidance values, specifically 0.1 and 0.2, in alignment with the parameters used in our proposed attack algorithm.
@@ 4.5 Ablation Studies
Figure 8: Comparison of the effectiveness of the objectness-only loss function versus the objectness $\times$ classification ($obj\times det$) loss function. The first row illustrates the results using the $obj$-only loss function, while the second row demonstrates the performance with the $obj\times det$ loss function. Empirical results consistently show that the $obj\times det$ approach outperforms the objectness-only method.

### v6

@@ Abstract
Adversarial attacks on deep learning models have received increased attention in recent years. Work in this area has mostly focused on gradient-based techniques, so-called “white-box” attacks, where the attacker has access to the targeted model’s internal parameters; such an assumption is usually untenable in the real world. Additionally, some attacks use the entire pixel space to fool a given model, which is neither practical nor physical. To accommodate these problems we propose the BBNP algorithm (Black-Box Naturalistic Patch): a direct, black-box, naturalistic, gradient-free method that uses the learned image manifold of a pretrained, generative adversarial network (GAN) to generate naturalistic adversarial patches for object detectors. This method performs model-agnostic black-box naturalistic attacks on object detection models by relying solely on the outputs of the model. Comparing our approach against five models, five black-box and two white-box attacks, we show that our proposed method achieves state-of-the-art results, outperforming all other tested black-box approaches.
@@ 3 Method
We exploit the smaller dimensionality of the GAN’s latent space by using an Evolution Strategy algorithm [30]. This method updates the latent vector z iteratively based on feedback from the fitness function. To compare, optimizing a patch directly in pixel space (of size $3\times 256\times 256$, as in our experiments) would involve $196,608$ pixels. In contrast, our approach operates in a $128$-dimensional latent space, achieving a reduction in dimensionality by a factor of 1,0001{,}000. This reduction makes the optimization process faster and the generated attacks more natural, thanks to the pretrained GAN, which is trained on ImageNet [31].
@@
Figure 1: Naturalistic Black-Box Adversarial Attack: Overview of framework. The system creates patches for object detectors by using the learned image manifold of a pretrained GAN generator ($G$) on real-world images. We use a pretrained classifier ($C$) to force the optimizer to find a patch that resembles a specific class, the TVTV component, which refers to total variation, aims to ensure the images are rendered with maximal smoothness, and the detector ($D$) for the actual detection loss. Efficient sampling of the GAN images via an iterative evolution strategy ultimately generates the final patch. The snowflake icon means frozen model weights.
@@ 3.2 Adversarial gradient estimation
$t$ is the timestep (epoch), $\alpha$ is the step size, and $\tilde{\nabla}_{z}\mathcal{L}_{total}$ is the gradient approximation of the total loss with respect to the latent vector $z$. We used $\tau=10$ in all experiments based on qualitative observations as follows. Figure 2 displays the patches generated without any form of projection. Conversely, Figure 5 presents the patches generated through the application of Equation 6.
@@ 4.1 Experimental setup
The experiments were performed using the Adam [38] optimizer, with a learning rate of $0.02$ and $\beta_{1}=0.5,\beta_{2}=0.999$. The optimization process involved populating the gradients through our evolutionary algorithm. If the changes in losses remain below 1e41e-4 for at least 50 epochs, we reduce the learning rate, following [10]. Our generator consists of BigGAN2 [27] (with frozen weights), which has been pretrained on the ImageNet dataset, with a latent vector of size $120$, and $128\times 128$ output resolution. As the BigGAN generator is class-conditional, the generated patch can be guaranteed to belong to a specific class. We used five different object detectors—Tiny-YOLOv3 [39], Tiny-YOLOv4 [40], YOLOv5s [41] [15, 42], SSD-L [43, 44] and L-DETR [45] trained on the COCO dataset [46]—with an input resolution of $416\times 416$. We chose these faster, memory-efficient models so as to be able to conduct more experiments. Moreover, these models represent architectures that are predominantly deployed on edge hardware and widely utilized in production environments [47, 48, 49], thereby enhancing the practical relevance and realism of the experiments conducted. For each of the detected bounding boxes, we placed an adversarial patch of size $25\%$ of the bounding box. We adopt the method proposed by Hu et al. [10] and generate a mask encompassing the clothing region for each individual, onto which we position our adversarial patch.
@@
We compared our proposed attack to five other black-box methods and two white-box methods: Square Attack [35], which exploits the CNN vulnerability to square perturbations; Sparse-RS [34], which leverages [35] to craft minimal modifications of squares; RS (pixel space random search) [34], which uses random search on pixel space; GAN Latent RS (latent-space random search) which uses random search in GAN latent space; and NES (natural evolution strategies) [36], which is an evolutionary attack performed in pixel space. We employed a consistent query budget of 110,000 for all attacks, derived from our chosen population size of 110 and 1,000 epochs.  [^1] We slightly adjusted some of the compared methods to accommodate the object-detection task: We removed the limitation on $\epsilon$ to encourage broader exploration during optimization in competitive attacks, because perturbation limits are not part of our targeted scenario, which is patch attacks. In our implementation of Square Attack, Sparse-RS, and the two white-box attacks, WB-Nat and WB-Adv as proposed by Thys et al. [6] and Hu et al. [10], we adhered strictly to the default parameters specified in their respective works. For NES, RS, and Latent-RS, we employed a noise standard deviation of $\sigma=0.1$. Additionally, for Latent-RS, we utilized two distinct classifier guidance values, specifically 0.1 and 0.2, in alignment with the parameters used in our proposed attack algorithm.
@@ 4.5 Ablation Studies
Figure 8: Comparison of the effectiveness of the objectness-only loss function versus the objectness $\times$ classification ($obj\times det$) loss function. The first row illustrates the results using the objobj-only loss function, while the second row demonstrates the performance with the $obj\times det$ loss function. Empirical results consistently show that the $obj\times det$ approach outperforms the objectness-only method.